ADNAN MUHAMMED

Grade 12,

Using identities, evaluate. 78 x 82 ??????????????????????

Using identities, evaluate.
78 x 82 ??????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

78 x 82
= (80 - 2)(80 + 2)
= 80^2- 2^2
= 6400 - 4
= 6396

Thanks